Short Report – QA/QC Meeting Aug 2021
Meeting Date: 2 Sep 2021
- From now on and starting next week, there will be semi-regular flux meetings for data questions.
- Database for our data is making good progress, there will be a meeting next week for next steps.
- Checking recent fluxes from our stations
LH, SB, MB, PM, TB, AG, IF, AS, LS, SE (10)
- If you are (Tech-)SRP, please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.
- SRP & Tech-SRP: please prepare short statement about your site and send to to LH (max. 2 sentences)
- The purpose of QA/QC meetings is to check on current, incoming data. SRPs choose specific issues we should look at together and discuss in the group. Fluxes are checked if the respective SRP wishes to do so.
- List of QA/QC Meeting dates: QA/QC Meetings 2021
CH-AWS is currently running OK, but there is a flux gap (approx. 2 weeks) in July. Please see GIN for details.
IF (SRP): CO2 and H2O flux, H and air temperature. Management: Mowing: 3.5, 12.6., 21.7. Herbicide: 13.8. Soil preparation and reseeding: 20.8. Slurry: 23.2., 7.5., 18.6., 29.7.
N2O flux and CH4 flux from Aerodyne QCL and LGR (from 22.7. onwards)
The LGR from CH-INO is now running at CH-CHA for N2O and CH4 fluxes, since 22 Jul 2021. There was grassland renewal (soil preparation and reseeding) on 20 Aug. High non-CO2 fluxes before and after renewal. High H in June coincided with high air temperatures.
No fluxes checked.
Last CO2, H2O, CH4 and N2O fluxes from INO@OE2 (station removed on 13.7.2021). Open-path IRGA -> more gaps. Three manure applications: 12.1. (solid), 4.3. (liquid) and 27.7. (liquid). Looks good.
Flux gap in May due to low IRGA flow rates.
RM (SRP): CO2 and H2O fluxes from 2021 over temporary grassland. Closed-path IRGA, no major gaps, already two cuts, third cut during today’s QA/QC meeting.